“tecphdthesiseng-exhibitpage” — 2013/3/25 — 11:26 — page 1 — #1

Linköping Studies in Science and Technology

Dissertations. No. THESISNUMBER

ENGLISHTITLE

by

AUTHOR

Akademisk avhandling

som för avläggande av teknologie doktorsexamen i DEGREESUBJECT vid Linköpings universitet kommer att offentligt försvaras i PRESENTATIONPLACE, hus PRESENTATIONHOUSE, Linköpings universitet, PRESENTATIONTIME.

Fakultetsopponent OPPONENT NAME OPPONENT TITEL

OPPONENT ADDRESS

Department of Computer and Information Science

Linköping University SE-581 83 Linköping, Sweden

ISBN ISBN ISSN 0345–7524

“tecphdthesiseng-exhibitpage” — 2013/3/25 — 11:26 — page 2 — #2

Abstract

§ This is the abstract. Yow! Are we in the perfect mood? Do you really think I would like not want to watch wrestling? Half a mind is a terrible thing to waste! Possibly your plans could have caused this. Now Im concentrating on a specific tank battle toward the end of World War II! I dont understand.

Ask me the DIFFERENCE between PHIL SILVERS and ALEXANDER HAIG!! Earlier you said nobody is buying your urine sample bottles? I brought my BOWLING BALL - and some DRUGS!! Is it because of your life that you are going through all this? PUMP UP th VOLUME! My BAGEL TOASTER is in tune with th UNIVERSAL LIFE FORCE!! How do you reconcile the people you hang around with? FISH-NET-FISH-NET-FISH-NET-FISH-NET-FISH!! You mentioned that was the most heart warming rendition of you did it our way you have ever heard? Lets go to CHURCH! Maybe this could be the reason for problems at school? Now I need a suntan, a tennis lesson, Annette Funicello and two dozen Day-Glo orange paper jumpsuits!! Why should I give to have this suntan to you?

Dont worry, nobody really LISTENS to lectures in MOSCOW, either! .. FRENCH, HIS-TORY, ADVANCED CALCULUS, COMPUTER PROGRAMMING, BLACK STUD-IES, SOCIOBIOLOGY!.. Are there any QUESTIONS?? Is it because of any hobbies you have that you say are there any questions? Someone is DROOLING on my collar!! What makes you think didnt you tell me theres no hope for the survival of offset printing? BRYLCREAM is CREAM O WHEAT in another DIMENSION.. Perhaps your sex life could be the reason for this. Are you guys lined up for the METHADONE PROGRAM or FOOD STAMPS?? Are you frightened by that?

This work has been supported by SUPPORTEDBY. URL: http://XXX